The NO molecule has a double degenerate electronic ground state and a double degenerate excited state at 121.1 cm- 1. Calculate and plot the electronic contribution to the molar heat capacity of the molecule up to 500K
Answer #1

Electronic partition function(qe) is given by:

qe=\sumgie-βEi

For NO , degeneracy , gi=2 ,β=1/kT , T=500K , Boltzmann constant ,k= 0.695cm-1/K

Transition in NO is from 2 Π1/2 -----> 2Π3/2

E1/2= 0 , E3/2 = 121.1cm-1

qe = 2 + 2 e-121.1/500*0.695 =2 + 2* 0.71 = 3.4